### v1

# Quantised helicity in optical media

Abstract
Optical helicity quantifies the handedness of light, and plays a central role in the description of interactions between light and chiral matter. In free space, it is related to the duality symmetry of the electromagnetic field, a continuous symmetry encapsulating the invariance of Maxwell’s equations under the interchange of electric and magnetic fields. However, in materials the situation is not so straightforward, as the free space transformation must be extended to encompass mixing of both the $\mathbf{E}$/$\mathbf{H}$ and $\mathbf{D}$/$\mathbf{B}$ field pairs. The simultaneous direct interchange of $\mathbf{E}$/$\mathbf{H}$ and of $\mathbf{D}$/$\mathbf{B}$ is incompatible with the presence of linear constitutive relations. In this work, we extend the duality transform in a way that resolves this incompatibility, and use this to define the optical helicity in a general medium, which may be dispersive, lossy, chiral or nonreciprocal. We find that the helicity density must contain an explicit contribution associated with the polarisation and magnetisation of the matter, and we show that the form of this matter contribution is independent of the details of the medium. We also show that the in-medium helicity can be naturally expressed in terms of the elementary quantised excitations of the system.
Helicity is an important and well-understood property in particle physics [1, 2], and in optics, where it is associated with circular polarisation [3]. The helicity is the fundamental measure of the handedness of light, and as such it has been identified as the key optical quantity in the differential action of light on chiral molecules and media [4, 5, 6]. This includes, in particular, chirally-selective forces present in a field with a helicity gradient [7, 8].
At low energies, that is in the absence of free charges and currents, Maxwell’s equations reduce to
$\boldsymbol{\nabla}\cdot\mathbf{D}$ $=0,$ $\boldsymbol{\nabla}\cdot\mathbf{B}$ $=0,$
×$\boldsymbol{\nabla}\times\mathbf{E}$ $=-\partial_{t}\mathbf{B},$$\boldsymbol{\nabla}\times\mathbf{H}$ $=\partial_{t}\mathbf{D}.$ (1)
It was Heaviside and Larmor [9, 10] who pointed out that these equations possess an electric-magnetic symmetry in that they are unchanged if we perform the rotations
$\mathbf{E}^{\prime}=\cos{\theta}\,\mathbf{E}+\sin{\theta}\,\mathbf{H},$
$\mathbf{H}^{\prime}=\cos{\theta}\,\mathbf{H}-\sin{\theta}\,\mathbf{E},$
$\mathbf{D}^{\prime}=\cos{\theta}\,\mathbf{D}+\sin{\theta}\,\mathbf{B},$
$\mathbf{B}^{\prime}=\cos{\theta}\,\mathbf{B}-\sin{\theta}\,\mathbf{D},$ (2)
where $\theta$ is a pseudoscalar, and we work in natural units in which $\varepsilon_{0}$, $\mu_{0}$, $c$ and $\hbar$ are all unity. In free space we can associate this symmetry with a locally conserved helicity density
$h_{\text{free}}=(\mathbf{A}\cdot\mathbf{B}-\mathbf{C}\cdot\mathbf{D})/2,$ (3)
where $\mathbf{A}$ is the magnetic vector potential, defined by $\mathbf{B}=\boldsymbol{\nabla}\times\mathbf{A}$, and $\mathbf{C}$ is the electric vector potential, defined by $\mathbf{D}=-\boldsymbol{\nabla}\times\mathbf{C}$ [11, 12]. The flux or flow of the helicity is simply the spin density [13, 14]. Indeed, in free space the Heaviside-Larmor transformation (Quantised helicity in optical media) may be interpreted as a rotation of the fields for each plane wave through an angle $\theta$ about its direction of propagation. As required the total helicity, which is the volume integral of $h_{\text{free}}$, acts as the generator of the transformation.
An important consequence of the Heaviside-Larmor symmetry is that all physical quantities must also be invariant under it [15]. This means, for example, that the energy density $(\mathbf{E}\cdot\mathbf{D}+\mathbf{B}\cdot\mathbf{H})/2$, and the kinetic and canonical momentum densities $\mathbf{E}\times\mathbf{H}/c^{2}$ and $\mathbf{D}\times\mathbf{B}$ [16], are satisfactory, but that $\varepsilon_{0}\mathbf{E}\times\mathbf{B}$ is not. This condition applies also to optical forces on a medium, and we note that the the Einstein-Laub force density [17]
$\mathbf{f}=(\mathbf{P}\cdot\boldsymbol{\nabla})\mathbf{E}+\frac{\partial\mathbf{P}}{\partial t}\times\mathbf{H}+(\mathbf{M}\cdot\boldsymbol{\nabla})\mathbf{H}-\frac{\partial\mathbf{M}}{\partial t}\times\mathbf{E},$ (4)
also satisfies the symmetry.
A problem arises when we seek the conserved quantity that is naturally associated with the Heaviside-Larmor transformation in a medium. If, for example, we consider the simplest constitutive relations, $\mathbf{D}=\varepsilon\mathbf{E}$ and $\mathbf{B}=\mu\mathbf{H}$, and impose the transformation from $\mathbf{E}$ and $\mathbf{H}$ to $\mathbf{E}^{\prime}$ and $\mathbf{H}^{\prime}$, then we find
$\mathbf{D}^{\prime}$ $=\varepsilon\mathbf{E}^{\prime}=\cos{\theta}\,\mathbf{D}+\sin{\theta}\,\left(\varepsilon/\mu\right)\mathbf{B},$
$\mathbf{B}^{\prime}$ $=\mu\mathbf{H}^{\prime}=\cos{\theta}\,\mathbf{B}-\sin{\theta}\,\left(\mu/\varepsilon\right)\mathbf{D},$ (5)
which is not the form required by the Heaviside-Larmor symmetry. This problem is depicted for both $\mathbf{D^{\prime}}$ and $\mathbf{B}^{\prime}$ in Figure 1. The transformations (Quantised helicity in optical media) and (Quantised helicity in optical media) are compatible only when $\varepsilon/\mu=1$. In this case, we say that the medium is dual-symmetric.
One way in which the incompatibility of (Quantised helicity in optical media), (Quantised helicity in optical media) and the constitutive relations has been addressed previously is by defining a duality transform in media which explicitly includes the permittivity and permeability [18, 19, 20, 21]. That is, to use the transformation
$\mathbf{E}^{\prime}$ $=\cos{\theta}\,\mathbf{E}+\sin{\theta}\,\sqrt{\mu/\varepsilon}\,\mathbf{H},$
$\mathbf{H}^{\prime}$ $=\cos{\theta}\,\mathbf{H}-\sin{\theta}\,\sqrt{\varepsilon/\mu}\,\mathbf{E},$
$\mathbf{D}^{\prime}$ $=\cos{\theta}\,\mathbf{D}+\sin{\theta}\,\sqrt{\varepsilon/\mu}\,\mathbf{B},$
$\mathbf{B}^{\prime}$ $=\cos{\theta}\,\mathbf{B}-\sin{\theta}\,\sqrt{\mu/\varepsilon}\,\mathbf{D},$ (6)
in place of (Quantised helicity in optical media). This transformation is a symmetry of Maxwell’s equations (1) under the condition that $\boldsymbol{\nabla}\sqrt{\varepsilon/\mu}=\mathbf{0}$ throughout space. This approach leads to an optical helicity density similar to the free space (3), with the addition of appropriate factors of $\varepsilon$ and $\mu$ [20]. In chiral media, the situation is more complicated, and the helicity density derived from the transformation above includes extra contributions proportional to the energy density in the medium [21].
Figure 1: Illustration comparing two ways of defining the duality transform in a medium. (a) The duality transform defined in equation (Quantised helicity in optical media) implies a simple rotation between the $\mathbf{D}$ and $\mathbf{B}$ fields; (b) The duality transforms defined in equations (Quantised helicity in optical media) and (Quantised helicity in optical media) are equivalent to a rotation and scaling, and alter both the relative lengths of $\mathbf{D}^{\prime}$ and $\mathbf{B^{\prime}}$ and the angle between them. It is clear that these transformations are both incompatible with those in equation (Quantised helicity in optical media).
The difference between the transformations (Quantised helicity in optical media) and (Quantised helicity in optical media) can be emphasised by noting that (Quantised helicity in optical media) corresponds to a rotation between $\mathbf{D}$ and $\mathbf{B}$, while (Quantised helicity in optical media) corresponds to a rotation and scaling, with a scale factor which depends on both $\varepsilon$ and $\mu$. Dispersion can be accommodated for in the transformation (Quantised helicity in optical media) [20]. From this point of view, treating dispersive media amounts to applying a different scaling for the different frequency components of the fields.
This approach works for linear, lossless materials, but there are features of the transformation (Quantised helicity in optical media) that make it very difficult to extend to more general cases. It is unclear how to account for frequency dependent absorption. Nonlinear responses are excluded from the outset, as the definition of the transformation (Quantised helicity in optical media) assumes that the field pairs $\mathbf{D}/\mathbf{B}$ and $\mathbf{E}/\mathbf{H}$ are connected by linear constitutive relations.
Furthermore, we may note that the transformations (Quantised helicity in optical media) imply the following transformation of $\mathbf{P}$ and $\mathbf{M}$,
$\mathbf{P}^{\prime}=\cos{\theta}\mathbf{P}+\sin\theta\sqrt{\frac{\varepsilon}{\mu}}\mathbf{M}+\sin\theta\Bigg{(}\sqrt{\frac{\varepsilon}{\mu}}-\sqrt{\frac{\mu}{\varepsilon}}\Bigg{)}\mathbf{H},$
$\mathbf{M}^{\prime}=\cos{\theta}\mathbf{M}-\sin{\theta}\sqrt{\frac{\mu}{\varepsilon}}\mathbf{P}+\sin\theta\Bigg{(}\sqrt{\frac{\varepsilon}{\mu}}-\sqrt{\frac{\mu}{\varepsilon}}\Bigg{)}\mathbf{E}.$ (7)
We therefore see that, while the transformations (Quantised helicity in optical media) may be a symmetry of Maxwell’s equations (1) even when $\varepsilon\neq\mu$, they then imply transformations of $\mathbf{P}$ and $\mathbf{M}$ that depend explicitly on the fields, rather than the simple rotation implied by (Quantised helicity in optical media).
All of these issues stem from the fact that (Quantised helicity in optical media) explicitly includes the constitutive parameters $\varepsilon$ and $\mu$ in the transformation. The approach of this Letter is to eschew (Quantised helicity in optical media), and to instead consider
$\mathbf{A}^{\prime}=\cos\theta\,\mathbf{A}+\sin\theta\,\mathbf{C},$ $\mathbf{C}^{\prime}=\cos\theta\,\mathbf{C}-\sin\theta\,\mathbf{A},$
$\mathbf{P}^{\prime}=\cos\theta\,\mathbf{P}+\sin\theta\,\mathbf{M},$ $\mathbf{M}^{\prime}=\cos\theta\,\mathbf{M}-\sin\theta\,\mathbf{P},$ (8)
as the duality transformation from which to derive the optical helicity inside a general medium. The key to this approach is that the fields $\mathbf{P}$ and $\mathbf{M}$ are not derived from the imposition of constitutive relations, but are explicitly included in the Lagrangian through an effective model of the material’s internal degrees of freedom (as in [22, 23]). The explicit inclusion of the material degrees of freedom in the Lagrangian leads to the natural emergence of a matter component in the medium helicity.
In order to enforce both of the transformations in Eq. (Quantised helicity in optical media) and, thereby, generate the Heaviside-Larmor transformation (Quantised helicity in optical media), we need to extend the form of the helicity and of its flux to include both field and matter components. This suggests that the helicity density should take the form
$h=\frac{1}{2}\left(\mathbf{A}\cdot\mathbf{B}-\mathbf{C}\cdot\mathbf{D}\right)+\boldsymbol{\Pi}^{P}\cdot\mathbf{M}-\boldsymbol{\Pi}^{M}\cdot\mathbf{P},\$ (9)
where $\boldsymbol{\Pi}^{P}$ and $\boldsymbol{\Pi}^{M}$ are the canonical momenta conjugate to $\mathbf{P}$ and $\mathbf{M}$. It is straightforward to confirm that the volume integral of this density generates the full Heaviside-Larmor transformation. If we promote our fields to quantum mechanical operators, then we find
$e^{i\theta\int d^{3}r\;\hat{h}(\mathbf{r})}\mathbf{\widehat{E}}(\mathbf{x})e^{-i\theta\int d^{3}r\;\hat{h}(\mathbf{r})}=\mathbf{\widehat{E}}(\mathbf{x})+\theta\mathbf{\widehat{H}}(\mathbf{x}),$ (10)
$e^{i\theta\int d^{3}r\;\hat{h}(\mathbf{r})}\mathbf{\widehat{H}}(\mathbf{x})e^{-i\theta\int d^{3}r\;\hat{h}(\mathbf{r})}=\mathbf{\widehat{H}}(\mathbf{x})-\theta\mathbf{\widehat{E}}(\mathbf{x}),$ (11)
$e^{i\theta\int d^{3}r\;\hat{h}(\mathbf{r})}\mathbf{\widehat{P}}(\mathbf{x})e^{-i\theta\int d^{3}r\;\hat{h}(\mathbf{r})}=\mathbf{\widehat{P}}(\mathbf{x})+\theta\mathbf{\widehat{M}}(\mathbf{x}),$ (12)
$e^{i\theta\int d^{3}r\;\hat{h}(\mathbf{r})}\mathbf{\widehat{M}}(\mathbf{x})e^{-i\theta\int d^{3}r\;\hat{h}(\mathbf{r})}=\mathbf{\widehat{M}}(\mathbf{x})-\theta\mathbf{\widehat{P}}(\mathbf{x}),$ (13)
to first order in $\theta$. The infinitesimal transformations of $\mathbf{D}$ and $\mathbf{B}$ follow directly. Finally, the flux density of our generalised helicity density is simply
$\mathbf{v}=\frac{1}{2}\left(\mathbf{E}\times\mathbf{A}+\mathbf{H}\times\mathbf{C}\right),$ (14)
which we recognise as the density of the electromagnetic spin.
For situations in which the helicity is locally conserved, the helicity density and this flux satisfy the local conservation law
$\frac{\partial h}{\partial t}+\boldsymbol{\nabla}\cdot\mathbf{v}=0,$ (15)
as may be derived directly from Noether’s theorem. The details of this are given in the Supplemental Material [24].
By explicitly including a matter component, we have obtained a generally applicable form for the helicity and established that it is indeed the conserved quantity associated with the full Heaviside-Larmor symmetry. It remains to examine how this quantity appears for a given medium and, moreover, to determine the physical forms of the two canonical momenta $\widehat{\mathbf{\Pi}}^{P}$ and $\widehat{\mathbf{\Pi}}^{M}$, which will depend on the mathematical model employed. To this end we introduce a model Lagrangian of a simple magnetodielectric medium.
In a magnetodielectric with a single electric and single magnetic resonance, the polarisation and magnetisation fields satisfy the equations of motion
$\frac{\partial^{2}\mathbf{P}}{\partial t^{2}}+\omega_{E}^{2}\mathbf{P}$ $=\alpha^{2}\mathbf{E},$
$\frac{\partial^{2}\mathbf{M}}{\partial t^{2}}+\omega_{M}^{2}\mathbf{M}$ $=\beta^{2}\mathbf{H},$ (16)
where $\omega_{E}$ and $\omega_{M}$ are the resonance frequencies, and $\alpha$ and $\beta$ are constants. With time-dependence $e^{-i\omega t}$, it is clear that these equations imply the constitutive relations $\mathbf{D}=\varepsilon\mathbf{E}$ and $\mathbf{B}=\mu\mathbf{H}$, with
$\varepsilon(\omega)=1+\frac{\alpha^{2}}{(\omega_{E}^{2}-\omega^{2})},$ $\mu(\omega)=1+\frac{\beta^{2}}{(\omega_{M}^{2}-\omega^{2})}.$ (17)
The equations of motion (16) are only invariant under the duality transform (Quantised helicity in optical media) if $\alpha=\beta$ and $\omega_{E}=\omega_{M}$. This means that $\varepsilon(\omega)=\mu(\omega)$ at all frequencies, and therefore that the electric and magnetic constituents of the medium emit in phase. The equations (16), along with Maxwell’s equations, may be obtained from the following Lagrangian density,
$\mathcal{L}=\frac{1}{2}[\mathbf{E}^{2}-\mathbf{B}^{2}]+\mathbf{P}\cdot\mathbf{E}+\mathbf{M}\cdot\mathbf{B}$ (18)
$+\frac{1}{2\alpha^{2}}\big{[}(\partial_{t}\mathbf{P})^{2}-\omega_{E}^{2}\mathbf{P}^{2}\big{]}+\frac{1}{2\beta^{2}}\big{[}(\partial_{t}\mathbf{M})^{2}-(\omega_{M}^{2}+\beta^{2})\mathbf{M}^{2}\big{]}.$
We take this Lagrangian as a starting point from which to derive an expression for the optical helicity. If we set $\alpha=\beta$ and $\omega_{E}=\omega_{M}=\omega_{0}$ then we can apply Noether’s theorem to recover the helicity directly [24]. Our Lagrangian provides the precise forms of the canonical momenta corresponding to the polarisation and magnetisation:
$\mathbf{\Pi}^{P}\equiv\frac{\partial\mathcal{L}}{\partial\dot{\mathbf{P}}}=\partial_{t}\mathbf{P}/\alpha^{2},$ (19)
$\mathbf{\Pi}^{M}\equiv\frac{\partial\mathcal{L}}{\partial\dot{\mathbf{M}}}=\partial_{t}\mathbf{M}/\beta^{2},$ (20)
and these may therefore be interpreted as bound polarisation and magnetisation currents for the Lagrangian (18). In the case of more general Lagrangians, this interpretation is not always possible. However, the expression for the helicity, when written in terms of the fields and their conjugate momenta, will still be given by equation (9).
We turn next to a microscopic quantum description of the helicity. In free space, the total helicity can be expressed as the difference between the number of left- and right-handed circularly polarised photons [25],
^$\widehat{\mathcal{H}}_{\text{free}}$ $=\int d^{3}r\,\hat{h}_{\text{free}}=\int d^{3}k\sum_{\lambda=\pm}\lambda\;\hat{a}^{\dagger}_{\lambda}(\mathbf{k})\hat{a}_{\lambda}(\mathbf{k}),$ (21)
where $\hat{a}_{\lambda}^{\dagger}(\mathbf{k})$ and $\hat{a}_{\lambda}(\mathbf{k})$ are the creation/annihilation operators for a photon of wavevector $\mathbf{k}$, and $\lambda=\pm 1$ labels the two circular polarisations. In a dual-symmetric medium, the helicity operator defined by equation (9) admits a similar interpretation: it is equal to the difference between the number of left- and right-handed polaritons in the system.
To see this, we diagonalise the system and examine the polariton modes. From the Lagrangian (18), we obtain the Hamiltonian
$\hat{H}=\int d^{3}k\;$ $\sum_{\lambda=\pm}\;k\;\hat{a}^{\dagger}_{\lambda}(\mathbf{k})\hat{a}_{\lambda}(\mathbf{k})$ (22)
$+\widetilde{\omega}_{0}\left[\hat{b}^{\dagger}_{\lambda}(\mathbf{k})\hat{b}_{\lambda}(\mathbf{k})+\hat{c}^{\dagger}_{\lambda}(\mathbf{k})\hat{c}_{\lambda}(\mathbf{k})\right]$
$-i\Lambda\left[\hat{a}_{\lambda}(\mathbf{k})-\hat{a}^{\dagger}_{\lambda}(-\mathbf{k})\right]\left[\hat{b}_{\lambda}(-\mathbf{k})+\hat{b}^{\dagger}_{\lambda}(\mathbf{k})\right]$
$-\lambda\Lambda\left[\hat{a}_{\lambda}(\mathbf{k})+\hat{a}^{\dagger}_{\lambda}(-\mathbf{k})\right]\left[\hat{c}_{\lambda}(-\mathbf{k})+\hat{c}^{\dagger}_{\lambda}(\mathbf{k})\right].$
where $\hat{a}_{\lambda}(\mathbf{k})$, $\hat{b}_{\lambda}(\mathbf{k})$ and $\hat{c}_{\lambda}(\mathbf{k})$ are the annihilation operators corresponding to the fields $\mathbf{A}$, $\mathbf{P}$ and $\mathbf{M}$. We have also defined $\widetilde{\omega}_{0}^{2}=\omega_{0}^{2}+\alpha^{2}$, as well as the coupling parameter $\Lambda=(\alpha/2)\sqrt{k/\widetilde{\omega}_{0}}$. We define the circular polarisation vectors $\mathbf{e}_{\pm}$ with the convention that that $\mathbf{e}_{\lambda}(\mathbf{k})\cdot\mathbf{e}_{\lambda^{\prime}}(\mathbf{k})=\delta_{\lambda\lambda^{\prime}}$, $\mathbf{e}_{\lambda}(\mathbf{k})\cdot\mathbf{k}=0$ for all $\lambda$, and $\mathbf{e}_{\lambda}(-\mathbf{k})=-\mathbf{e}_{-\lambda}(\mathbf{k})$. Our continuum annihilation and creation operators satisfy the bosonic commutation relations $\left[\hat{a}_{\lambda}(\mathbf{k}),\hat{a}^{\dagger}_{\lambda^{\prime}}(\mathbf{k}^{\prime})\right]=\delta_{\lambda\lambda^{\prime}}\delta(\mathbf{k}-\mathbf{k}^{\prime})$, with similar relations for $\hat{b}_{\lambda}(\mathbf{k})$ and $\hat{c}_{\lambda}(\mathbf{k})$, and all other commutators equal to zero. The explicit expansions of $\hat{\mathbf{A}}$, $\hat{\mathbf{P}}$ and $\hat{\mathbf{M}}$ used to obtain (22) from (18) are given in the Supplemental Material [24].
To diagonalise (22), we follow [22] and define polariton annihilation operators as linear combinations of the $\mathbf{A}$, $\mathbf{P}$ and $\mathbf{M}$ creation/annihilation operators,
$\hat{p}_{i}(\mathbf{k},\lambda)=\alpha_{i}\hat{a}_{\lambda}(\mathbf{k})$ $+\widetilde{\alpha}_{i}\hat{a}^{\dagger}_{\lambda}(-\mathbf{k})+\beta_{i}\hat{b}_{\lambda}(\mathbf{k})+\widetilde{\beta}_{i}\hat{b}^{\dagger}_{\lambda}(-\mathbf{k})$
$+\gamma_{i}\hat{c}_{\lambda}(\mathbf{k})+\widetilde{\gamma}_{i}\hat{c}^{\dagger}_{\lambda}(-\mathbf{k}),$ (23)
with the requirement that $\left[\hat{p}_{i}(\mathbf{k},\lambda),\hat{H}\right]=\omega_{i}\hat{p}_{i}(\mathbf{k},\lambda)$. The coefficients $\alpha_{i}$, $\widetilde{\alpha}_{i}$, $\beta_{i}$, $\widetilde{\beta}_{i}$, $\gamma_{i}$ and $\widetilde{\gamma}_{i}$ are constants, and the index $i$ labels the different polariton branches. The details of the diagonalisation are given in the supplemental material [24]. We find that the polariton branch frequencies, $\omega_{i}$, satisfy the dispersion relation
$k^{2}=\omega_{i}^{2}n^{2}(\omega_{i})=\omega_{i}^{2}\left(1-\frac{\alpha^{2}}{\omega_{i}^{2}-\omega_{0}^{2}}\right)^{2}$ (24)
the solutions of which are plotted in Fig. 2.
Figure 2: The polariton branches, i.e. solutions to the dispersion relation in Eq. (24), with $\alpha=\omega_{0}/2$, chosen to ensure that the wave-vector and group velocity direction coincide.
Once we have the explicit expressions for the polariton operators, we are in a position to express the helicity in terms of the polaritons. After some algebra, we obtain
$:\widehat{\mathcal{H}}:$ $=\int d^{3}r\;:\frac{1}{2}\left(\mathbf{\widehat{A}}\cdot\mathbf{\widehat{B}}-\mathbf{\widehat{C}}\cdot\mathbf{\widehat{D}}\right)+\boldsymbol{\widehat{\Pi}}^{P}\cdot\mathbf{\widehat{M}}-\boldsymbol{\widehat{\Pi}}^{M}\cdot\mathbf{\widehat{P}}:$
$=\sum_{\lambda=\pm}\int d^{3}k\;\lambda\;\bigg{[}\hat{p}^{\dagger}_{u}(\mathbf{k},\lambda)\hat{p}_{u}(\mathbf{k},\lambda)+\hat{p}^{\dagger}_{d}(\mathbf{k},\lambda)\hat{p}_{d}(\mathbf{k},\lambda)$
$\hskip 85.35826pt-\hat{p}^{\dagger}_{p}(\mathbf{k},\lambda)\hat{p}_{p}(\mathbf{k},\lambda)\bigg{]},$ (25)
where we have normal-ordered with respect to the polariton vacuum, and have labelled the three polariton branches $u$, $d$ and $p$. Equation (Quantised helicity in optical media) is simply the difference between the number of positive and negative helicity polaritons, and is completely analogous to the free space expression (21). Note that one of the modes has a negative refractive index, and therefore contributes to the total helicity with opposite sign (as for this mode, the wave propagates in the opposite direction to the wave vector).
We note that the helicity is only expressible in this simple way when the medium is dual-symmetric (i.e. when $\alpha=\beta$ and $\omega_{E}=\omega_{M}$). This is to be expected as the helicity is only expressible in terms of polariton number operators in a dual-symmetric medium, as in a general lossless (but non-dual-symmetric) medium, the number of polaritons is conserved while the helicity is not. In such a medium, the total helicity operator can still be expressed in terms of circularly polarised polaritons, but will include more general products of polariton creation and annihilation operators.
Discussion. In this Letter, we set out to find an expression for the optical helicity which is valid within any optical medium, and to resolve the mutual incompatibility between the duality transform (Quantised helicity in optical media), and the constitutive equations. We approached this by using the duality transform (Quantised helicity in optical media) as the starting point from which to define the optical helicity, which provides the appropriate rotation between $\mathbf{E}$ and $\mathbf{H}$, and also $\mathbf{D}$ and $\mathbf{B}$, without coming into conflict with constitutive relations or presupposing any particular properties of the medium. This works because, in our approach, the constitutive equations connecting the fields emerge as a consequence of the Lagrangian and the associated equations of motion, rather than being enforced from the outset. We therefore require a description of the material’s internal degrees of freedom in the Lagrangian, and this explicit description of the material causes a medium contribution to the helicity to naturally emerge.
The extension of the Lagrangian (18) to allow for multiple resonances in $\varepsilon$ and $\mu$ may be achieved by straightforward modification of the harmonic oscillator terms. Lossy media can be treated by coupling the $\mathbf{P}$ and $\mathbf{M}$ fields to a continuous reservoir of harmonic oscillators [26, 27, 28, 29, 30]. Finally, chiral and magnetoelectric responses can also be added to these models by introducing additional terms into $\mathcal{L}$ which couple the $\mathbf{P}$ and $\mathbf{M}$ degrees of freedom [31]. As we have argued, any such modifications will not affect the basic results (9) and (14).
When comparing our results to those obtained elsewhere in the literature, it is important to appreciate that (Quantised helicity in optical media) and (Quantised helicity in optical media) are, in general, distinct symmetry transformations, and so one would not expect the conserved quantities associated with these to coincide exactly. Even in the case where $\varepsilon=\mu$, in which (Quantised helicity in optical media) reduces to (Quantised helicity in optical media), we note that the helicity arising from Eq. (Quantised helicity in optical media) contains a matter contribution absent from the helicity derived from Eq. (Quantised helicity in optical media). This situation is analogous to that discussed in Ref. [23] which points out that the total electromagnetic energy used in phenomenological quantisation schemes differs from the total energy of the full system by a matter component.
A point worth mentioning is that the definition presented here relies on the ability to use a model for the medium that can be made symmetrical under the duality transform (Quantised helicity in optical media), for at least some choice of model parameters. This requires that the polarisation and magnetisation can be written as functions of $\mathbf{D}$ and $\mathbf{B}$, or $\mathbf{E}$ and $\mathbf{H}$, but not some combination of the two. For example, a medium where $\mathbf{P}=\mathbf{P}(\mathbf{E},\mathbf{H})$ and $\mathbf{M}=\mathbf{M}(\mathbf{H},\mathbf{E})$ would fulfil this requirement, but one where $\mathbf{P}=\mathbf{P}(\mathbf{E},\mathbf{B})$ and $\mathbf{M}=\mathbf{M}(\mathbf{B},\mathbf{E})$ would not. It is possible to reformulate the description of such materials in a way that can be made symmetric under the duality transform (Quantised helicity in optical media); we leave this to future work.
As a final point, we may note that, unlike the helicity density $h$, the helicity flux density, $\mathbf{v}$, derived from our procedures contains no additional medium component. This makes physical sense, as the medium constituents cannot themselves flow, leaving the task of carrying helicity from place to place entirely to the electromagnetic fields.
Acknowledgements.
Acknowledgements. This research was supported by the Royal Commission for the Exhibition of 1851, the Royal Society (grant numbers RSRP/RE/220013 and RP/150122), the EPSRC (grant number EP/V048449/1) and the Leverhulme Trust.
Supplemental Information for
Quantised helicity in optical media
I Noether’s Theorem
Here we provide the details of the application of Noether’s theorem [32] to derive an expression for the optical helicity from the Lagrangian (18). Under the infinitesimal version of duality transformation (Quantised helicity in optical media),
$\mathbf{A}^{\prime}=\mathbf{A}+\delta\theta\mathbf{C},$ $\mathbf{C}^{\prime}=\mathbf{C}-\delta\theta\mathbf{A},$
$\mathbf{P}^{\prime}=\mathbf{P}+\delta\theta\mathbf{M},$ $\mathbf{M}^{\prime}=\mathbf{M}-\delta\theta\mathbf{P},$ (SI 1)
the Lagrangian becomes $\mathcal{L}^{\prime}=\mathcal{L}(\mathbf{A^{\prime}},\mathbf{C^{\prime}},\mathbf{P^{\prime}},\mathbf{M^{\prime}})$. We first consider the change in the Lagrangian, $\delta\mathcal{L}\equiv\mathcal{L}^{\prime}-\mathcal{L}$. Substituting the duality transformation into the Lagrangian (18), we find that
$\delta\mathcal{L}$ $=\delta\theta\left(\mathbf{E}\cdot\mathbf{B}+\mathbf{H}\cdot\mathbf{D}\right)$ (SI 2)
$=\frac{\delta\theta}{2}\left(-\partial_{t}\left[\mathbf{A}\cdot\mathbf{B}+\mathbf{C}\cdot\mathbf{D}\right]-\boldsymbol{\nabla}\cdot\left[\mathbf{E}\times\mathbf{A}-\mathbf{H}\times\mathbf{C}\right]\right),$
which is a four-divergence, showing that the equations of motion remain unchanged under the transform [33, 34].
[…]
where we have used that $\boldsymbol{\Pi}^{P}=\partial_{t}\mathbf{P}/\alpha^{2}$ and $\boldsymbol{\Pi}^{M}=\partial_{t}\mathbf{M}/\alpha^{2}$ in the chosen gauge and coupling. Equating the two expressions for $\delta\mathcal{L}$ from Eq. (SI 2) and Eq. (SI 3) leads to the local conservation law
$\frac{\partial}{\partial t}\bigg{[}\frac{1}{2}\bigg{(}\mathbf{A}\cdot\mathbf{B}$ $-\mathbf{C}\cdot\mathbf{D}\bigg{)}+\boldsymbol{\Pi}^{P}\cdot\mathbf{M}-\boldsymbol{\Pi}^{M}\cdot\mathbf{P}\bigg{]}$
$+\boldsymbol{\nabla}\cdot\left[\frac{1}{2}\left(\mathbf{E}\times\mathbf{A}+\mathbf{H}\times\mathbf{C}\right)\right]=0,$ (SI 4)
which explicitly shows the helicity density (9) and flux (14) from the main text.
II Hamiltonian in field variables
If we consider a dual-symmetric medium (that is, we set $\alpha=\beta$ and $\omega_{E}=\omega_{M}=\omega_{0}$), then the Hamiltonian obtained from the Lagrangian in Eq. (18) takes the form
$\hat{H}$ $=\int d^{3}r\;\frac{1}{2}\left[\left(\boldsymbol{\widehat{\Pi}}^{A}\right)^{2}+\mathbf{\widehat{B}}^{2}\right]$
$\hskip 42.67912pt+\frac{1}{2\alpha^{2}}\left[\alpha^{4}\left(\boldsymbol{\widehat{\Pi}}^{P}\right)^{2}+\left(\omega_{0}^{2}+\alpha^{2}\right)\mathbf{\widehat{P}}^{2}\right]$
$\hskip 42.67912pt+\frac{1}{2\alpha^{2}}\left[\alpha^{4}\left(\boldsymbol{\widehat{\Pi}}^{M}\right)^{2}+\left(\omega_{0}^{2}+\alpha^{2}\right)\mathbf{\widehat{M}}^{2}\right]$
$\hskip 42.67912pt+\mathbf{\widehat{P}}\cdot\boldsymbol{\widehat{\Pi}}^{A}-\mathbf{\widehat{M}}\cdot\mathbf{\widehat{B}},$
which is dual symmetric with respect to the transform in $(\mathbf{\widehat{D}},\mathbf{\widehat{B}})$, as $\boldsymbol{\widehat{\Pi}}^{A}=-\mathbf{\widehat{D}}$. If we expand the fields as
$\mathbf{\widehat{A}}$ $=\sum_{\lambda=\pm}\int\frac{d^{3}k}{(2\pi)^{3/2}}\;\frac{1}{\sqrt{2k}}\bigg{[}\mathbf{e}_{\lambda}(\mathbf{k})\hat{a}_{\lambda}(\mathbf{k})e^{i\mathbf{k}\cdot\mathbf{x}}+\mathbf{e}_{-\lambda}(\mathbf{k})\hat{a}^{\dagger}_{\lambda}(\mathbf{k})e^{-i\mathbf{k}\cdot\mathbf{x}}\bigg{]},$
$\mathbf{\widehat{P}}$ $=\sum_{\lambda=\pm}\int\frac{d^{3}k}{(2\pi)^{3/2}}\;\frac{\alpha}{\sqrt{2\widetilde{\omega}_{0}}}\bigg{[}\mathbf{e}_{\lambda}(\mathbf{k})\hat{b}_{\lambda}(\mathbf{k})e^{i\mathbf{k}\cdot\mathbf{x}}+\mathbf{e}_{-\lambda}(\mathbf{k})\hat{b}^{\dagger}_{\lambda}(\mathbf{k})e^{-i\mathbf{k}\cdot\mathbf{x}}\bigg{]},$ (SI 5)
$\mathbf{\widehat{M}}$ $=\sum_{\lambda=\pm}\int\frac{d^{3}k}{(2\pi)^{3/2}}\;\frac{\alpha}{\sqrt{2\widetilde{\omega}_{0}}}\bigg{[}\mathbf{e}_{\lambda}(\mathbf{k})\hat{c}_{\lambda}(\mathbf{k})e^{i\mathbf{k}\cdot\mathbf{x}}+\mathbf{e}_{-\lambda}(\mathbf{k})\hat{c}^{\dagger}_{\lambda}(\mathbf{k})e^{-i\mathbf{k}\cdot\mathbf{x}}\bigg{]},$
then we arrive at the Hamiltonian seen in the main text (22).
III Details of the diagonalisation
Here we present the details of the diagonalisation of (22), leading to expressions for the Hopfield coefficients appearing in (Quantised helicity in optical media). We begin with some preparatory results, which will be useful in expanding the helicity operator to obtain Eq. (Quantised helicity in optical media) . First, we can invert Eq. (Quantised helicity in optical media) by noting that, as the polariton operators form a complete basis set, we must have that $\hat{a}_{\lambda}(\mathbf{k})=\sum_{i}f_{i}\hat{p}_{i}(\mathbf{k},\lambda)+g_{i}\hat{p}^{\dagger}_{i}(-\mathbf{k},\lambda)$ for some constants $f_{i}$ and $g_{i}$ (and similarly for the other operators). Computing the commutators $\left[\hat{a}_{\lambda}(\mathbf{k}^{\prime}),\hat{p}_{i}(\mathbf{k},\lambda)\right]$ and $\left[\hat{a}_{\lambda}(\mathbf{k}^{\prime}),\hat{p}^{\dagger}_{i}(-\mathbf{k},\lambda)\right]$ allows us to identify the constants,
$\hat{a}_{\lambda}(\mathbf{k})$ $=\sum_{i}\alpha_{i}^{*}\hat{p}_{i}(\mathbf{k},\lambda)-\widetilde{\alpha}_{i}\hat{p}^{\dagger}_{i}(-\mathbf{k},\lambda),$
$\hat{b}_{\lambda}(\mathbf{k})$ $=\sum_{i}\beta_{i}^{*}\hat{p}_{i}(\mathbf{k},\lambda)-\widetilde{\beta}_{i}\hat{p}^{\dagger}_{i}(-\mathbf{k},\lambda),$ (SI 6)
$\hat{c}_{\lambda}(\mathbf{k})$ $=\sum_{i}\gamma_{i}^{*}\hat{p}_{i}(\mathbf{k},\lambda)-\widetilde{\gamma}_{i}\hat{p}^{\dagger}_{i}(-\mathbf{k},\lambda),$
Furthermore, inserting the expressions in Eqns. (SI 6) into the commutators $\left[\hat{a}_{\lambda}(\mathbf{k}),\hat{a}^{\dagger}_{\lambda^{\prime}}(\mathbf{k}^{\prime})\right]=\delta_{\lambda\lambda^{\prime}}\delta(\mathbf{k}-\mathbf{k}^{\prime})$, etc., produces the relations
$\sum_{i}|\alpha_{i}|^{2}-|\widetilde{\alpha}_{i}|^{2}$ $=1,$
$\sum_{i}|\beta_{i}|^{2}-|\widetilde{\beta}_{i}|^{2}$ $=1,$ (SI 7)
$\sum_{i}|\gamma_{i}|^{2}-|\widetilde{\gamma}_{i}|^{2}$ $=1,$
which are types of sum rules. These sum rules are important, and we must always make sure that they are respected in order to have proper commutators.
We now derive the explicit expressions for the Hopfield coefficients. Using $\left[\hat{p}_{i}(\mathbf{k},\lambda),\hat{H}\right]=\omega_{i}\hat{p}_{i}(\mathbf{k},\lambda)$ [22], along with the commutators for the creation and annihilation operators, implies the algebraic relations
$(\omega_{i}-k)\alpha_{i}$ $=\Lambda\left[-i(\beta_{i}-\widetilde{\beta}_{i})-\lambda(\gamma_{i}-\widetilde{\gamma}_{i})\right],$
$(\omega_{i}+k)\widetilde{\alpha}_{i}$ $=\Lambda\left[i(\beta_{i}-\widetilde{\beta}_{i})-\lambda(\gamma_{i}-\widetilde{\gamma}_{i})\right],$
$\beta_{i}$ $=i\Lambda\left(\frac{\alpha_{i}+\widetilde{\alpha}_{i}}{\omega_{i}-\widetilde{\omega}_{0}}\right),$
$\widetilde{\beta}_{i}$ $=i\Lambda\left(\frac{\alpha_{i}+\widetilde{\alpha}_{i}}{\omega_{i}+\widetilde{\omega}_{0}}\right),$ (SI 8)
$\gamma_{i}$ $=-\lambda\Lambda\left(\frac{\alpha_{i}-\widetilde{\alpha}_{i}}{\omega_{i}-\widetilde{\omega}_{0}}\right),$
$\widetilde{\gamma}_{i}$ $=-\lambda\Lambda\left(\frac{\alpha_{i}-\widetilde{\alpha}_{i}}{\omega_{i}+\widetilde{\omega}_{0}}\right).$
Along with this, we also get a normalisation from the the commutator $\left[\hat{p}_{i}(\mathbf{k},\lambda),\hat{p}^{\dagger}_{j}(\mathbf{k}^{\prime},\lambda^{\prime})\right]=\delta_{ij}\delta_{\lambda\lambda^{\prime}}\delta(\mathbf{k}-\mathbf{k}^{\prime})$:
$|\alpha_{i}|^{2}-|\widetilde{\alpha}_{i}|^{2}+|\beta_{i}|^{2}-|\widetilde{\beta}_{i}|^{2}+|\gamma_{i}|^{2}-|\widetilde{\gamma}_{i}|^{2}=1.$ (SI 9)
This completely solves the system, and leads to explicit expressions for our coefficients $\alpha_{i}$, $\widetilde{\alpha}_{i}$, $\beta_{i}$, $\widetilde{\beta}_{i}$, $\gamma_{i}$ and $\widetilde{\gamma}_{i}$. It is important to note, however, that one of the normal mode frequencies defined through Eq. (24) has a negative group velocity. In fact, if we simplify Eq. (24) and pick the positive part of the square root such that
$n(\omega_{i})=\frac{k}{\omega_{i}}=\left(1-\frac{\alpha^{2}}{\omega_{i}^{2}-\omega_{0}^{2}}\right),$ (SI 10)
then we find two positive frequency modes and one negative. If we on the other hand choose the negative square root, we arrive at
$n(\omega_{i})=\frac{k}{\omega_{i}}=-\left(1-\frac{\alpha^{2}}{\omega_{i}^{2}-\omega_{0}^{2}}\right),$ (SI 11)
along with two negative frequency and one positive frequency mode. We want to choose the positive frequency (positive norm) modes, and notation simplifies if we define
$n_{u/d}(\omega_{u/d})=1-\frac{\alpha^{2}}{\omega_{u/d}^{2}-\omega_{0}^{2}};\hskip 2.84544ptn_{p}(\omega_{p})=-1+\frac{\alpha^{2}}{\omega_{p}^{2}-\omega_{0}^{2}},$
where we have chosen the modes $\omega_{i}=\{\omega_{u},\omega_{p},\omega_{d}\}$ in descending order of oscillation frequency. Manipulating the algebratic relations in Eqns. (III) yields
$\left(\frac{\omega_{i}}{k}-1\right)\alpha_{i}+\left(\frac{\omega_{i}}{k}+1\right)\widetilde{\alpha}_{i}=(\alpha_{i}-\widetilde{\alpha}_{i})\frac{\alpha^{2}}{\omega_{i}^{2}-\widetilde{\omega}_{0}^{2}},$
at which stage we must distinguish between $\omega_{u/d}$ and $\omega_{p}$. For $\omega_{u/d}$ we have
$\left(\frac{\omega_{u/d}}{k}-1\right)\alpha_{u/d}$ $+\left(\frac{\omega_{u/d}}{k}+1\right)\widetilde{\alpha}_{u/d}$
$=\left(\alpha_{u/d}-\widetilde{\alpha}_{u/d}\right)\left(\frac{\omega_{u/d}}{k}-1\right)$
$\Rightarrow\widetilde{\alpha}_{u/d}=0,$
whereas for $\omega_{p}$ we find that
$\left(\frac{\omega_{p}}{k}-1\right)\alpha_{p}$ $+\left(\frac{\omega_{p}}{k}+1\right)\widetilde{\alpha}_{p}$
$=(-\alpha_{p}+\widetilde{\alpha}_{p})\left(\frac{\omega_{p}}{k}+1\right)$
$\Rightarrow\alpha_{p}=0.$
It is now straightforward to show that
$\left|\alpha_{u/d}\right|^{2}$ $=\left(1+2\Lambda^{2}\left[\frac{1}{(\omega_{u/d}-\widetilde{\omega}_{0})^{2}}-\frac{1}{(\omega_{u/d}+\widetilde{\omega}_{0})^{2}}\right]\right)^{-1}$ (SI 12)
$\left|\widetilde{\alpha}_{p}\right|^{2}$ $=\left(-1+2\Lambda^{2}\left[\frac{1}{(\omega_{p}-\widetilde{\omega}_{0})^{2}}-\frac{1}{(\omega_{p}+\widetilde{\omega}_{0})^{2}}\right]\right)^{-1}$ (SI 13)
using Eqns. (III) and the normalisation condition Eq. (SI 9). Importantly, we can note that this satisfies
$\left|\alpha_{u}\right|^{2}+\left|\alpha_{d}\right|^{2}-\left|\widetilde{\alpha}_{p}\right|^{2}=1,$
which must be the case in order to preserve the canonical commutation relations. The phases of $\alpha_{u/d}$ and $\widetilde{\alpha}_{p}$ can be chosen freely, and the remaining Hopfield coefficients follow from Eqns. (III).
IV Diagonal form of the helicity
In this section, we provide details of how to obtain Eq. (Quantised helicity in optical media) in the main text, using the polariton coefficients obtained in the previous section. Let us express the total helicity, Eq. (9), in terms of the photon and matter annihilation and creation operators, yielding:
$\widehat{\mathcal{H}}$ $=\int d^{3}r\;\left[\frac{1}{2}\left(\mathbf{\widehat{A}}\cdot\mathbf{\widehat{B}}-\mathbf{\widehat{C}}\cdot\mathbf{\widehat{D}}\right)+\boldsymbol{\widehat{\Pi}}^{P}\cdot\mathbf{\widehat{M}}-\boldsymbol{\widehat{\Pi}}^{M}\cdot\mathbf{\widehat{P}}\right]$
$=\sum_{\lambda=\pm}\int d^{3}k\;\lambda\;\hat{a}^{\dagger}_{\lambda}(\mathbf{k})\hat{a}_{\lambda}(\mathbf{k})$
$\hskip 56.9055pt+i\left(\hat{c}_{\lambda}(\mathbf{k})\hat{b}^{\dagger}_{\lambda}(\mathbf{k})-\hat{c}^{\dagger}_{\lambda}(\mathbf{k})\hat{b}_{\lambda}(\mathbf{k})\right).$
This can further be rewritten in terms of polaritons using Eqns. (SI 6), where we should note that we need to normal order the operators with respect to the polariton vacuum. This yields
$:\hat{h}_{\text{total}}:=\sum_{\lambda=\pm}$ $\int d^{3}k\;\sum_{ij}\bigg{(}\lambda\left(\alpha_{i}\alpha_{j}^{*}+\widetilde{\alpha}_{i}\widetilde{\alpha}_{j}^{*}\right)$
$+2\text{Im}\left[\beta_{i}^{*}\gamma_{j}+\widetilde{\gamma}_{i}^{*}\widetilde{\beta}_{j}\right]\bigg{)}\hat{p}^{\dagger}_{i}(\mathbf{k},\lambda)\hat{p}_{j}(\mathbf{k},\lambda)$
+$+\bigg{\{}\left(2\text{Im}\left[\widetilde{\gamma}_{i}\beta_{j}-\gamma_{i}\widetilde{\beta}_{j}\right]-\lambda\alpha_{i}\widetilde{\alpha}_{j}\right)$
[…]
We can simplify this by using Eq. (III) to express $\beta_{i}$, $\widetilde{\beta}_{i}$, $\gamma_{i}$ and $\widetilde{\gamma}_{i}$ in terms of $\alpha_{i}$ and $\widetilde{\alpha}_{i}$. We may then use the fact that $\widetilde{\alpha}_{u/d}=0$ and $\alpha_{p}=0$, as well as the explicit solutions for $\alpha_{u/d}$ and $\widetilde{\alpha}_{p}$ given in Eq. (SI 12) and Eq. (SI 13), to show that the $i\neq j$ terms vanish.
In order to show this cancellation, we also must make use of the fact that when $i\neq j$, the frequencies $\omega_{i}$ and $\omega_{j}$ are related to one another, as both are two distinct roots of the same dispersion relation. The relevant dispersion relations Eqns. (SI 10) and  (SI 11) are cubics in $\omega$, so sums and products of the two roots which appear can be expressed in terms of the third root. Finally, using Eq. (SI 12) and Eq. (SI 13) again, we find Eq. (Quantised helicity in optical media) from the main text.